18 The treasurer of Riley Coal Co. is asked to compute the cost of fixed income...

18

The treasurer of Riley Coal Co. is asked to compute the cost of fixed income securities for her corporation. Even before making the calculations, she assumes the aftertax cost of debt is at least 4 percent less than that for preferred stock.

Debt can be issued at a yield of 12.0 percent, and the corporate tax rate is 25 percent. Preferred stock will be priced at $62 and pay a dividend of $7.40. The flotation cost on the preferred stock is $7.


a. Compute the aftertax cost of debt. (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)


b. Compute the aftertax cost of preferred stock. (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

c. Based on the facts given above, is the treasurer correct?
  

No, the treasurer is incorrect.
Yes, the treasurer is correct.
Answer #1

a

After tax rate = YTM * (1-Tax rate)
After tax rate = 12 * (1-0.25)
After tax rate = 9

b

Cost of preferred equity= dividend/(Price-floatation charge) = 7.4/(62-7)=13.45%

c

Difference = 13.45-9=4.45%